A 1305 kg test car travels around a flat circular track that is 2.15 km long (circumference). If the coefficient of friction between the car’s tires and the track is 1.20 will the car be able to drive a speed of 55.7 m/s and stay on the track?



1
Expert's answer
2021-12-13T16:56:04-0500

The Newton's second law says

"ma=F=\mu N=\mu mg"

"v^2/R=\mu g"

The critical speed

"v_c=\sqrt{\mu gR}=\sqrt{1.20*9.8*2150}=159\:\rm m/s"

Because "v<v_c" , so the car be able to drive a speed of 55.7 m/s and stay on the track.